# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v1

@@ Abstract
Electron spin-qubits in silicon-germanium (SiGe) heterostructures are a major candidate for the realization of scalable quantum computers. A critical challenge in strained Si/SiGe quantum wells (QWs) is the existence of two nearly degenerate valley states at the conduction band minimum that can lead to leakage of quantum information. To address this issue, various strategies have been explored to enhance the valley splitting (i.e., the energy gap between the two low-energy conduction band minima), such as sharp interfaces, oscillating germanium concentrations in the QW and shear strain engineering. In this work, we develop a comprehensive envelope-function theory augmented by the empirical pseudopotential method to incorporate the effects of alloy disorder, strain, and non-trivial resonances arising from interaction between valley states across different Brillouin zones. We apply our model to analyze common epitaxial profiles studied in the literature and compare our results with previous work. This framework provides an efficient tool for quantifying the interplay of these effects on the valley splitting, enabling complex epitaxial profile optimization in future work.
@@ I Introduction
Electron spin-qubits in Si/SiGe heterostructures are one of the major candidates for the realization of fault-tolerant universal quantum computers [1, 2, 3]. The material platform has excellent scalability prospects because of the abundance of nuclear spin free isotopes (e.g., 28Si and 76Ge) required for long coherence times and its compatibility with industrial fabrication technology [4, 5, 6]. Experiments have demonstrated high-fidelity state initialization and readout in combination with one and two-qubit gates exceeding the fault-tolerance threshold [7, 8, 9]. In addition, scalable quantum computing architectures require coherent coupling of distant qubits to overcome crosstalk and qubit wiring limitations [10, 11]. As a major step in this direction, coherent qubit transfer across the chip was recently demonstrated using conveyor-mode electron spin-qubit shuttles [12, 13].
@@
One of the key challenges in the design of reliable Si/SiGe qubits is the enhancement of the energy splitting between the two nearly degenerate valley states at the conduction band minimum of a biaxially strained SiGe/Si/SiGe quantum well (QW), see Fig. 1. The energy splitting between these states, called valley splitting, is caused by the coupling of the two valley states by the heterostructure potential [14, 15, 16]. Interface roughness and random alloy disorder in the SiGe barrier [17, 18, 19] cause the valley splitting to fluctuate across the chip with typical values ranging from several tens to hundreds of µeV [20, 21, 22, 23]. These statistical fluctuations of the valley splitting are notoriously difficult to control and inevitably lead to spatial domains with very low splitting, where it becomes comparable with the Zeeman-splitting [24, 25, 22]. As a result, low valley splittings lead to so-called spin-valley hotspots, which are a potential source for spin-dephasing and leakage of quantum information [26, 27]. While these issues might be secondary for stationary qubits, they are particularly challenging in spin-qubit shuttles, where the electron is conveyed over micrometer distances across a disordered landscape [28, 29]. For such applications a reliably large valley splitting above 200 µeV would be desirable to avoid spin-valley hotspots [30, 28, 31].
@@
Fig. 1: (a) First Brillouin zone of the face-centered cubic (fcc) lattice. The degeneracy between the six equivalent conduction band minima in Si near the $X$-points is lifted by strain. (b) Energy diagram for the conduction band ground states in Si/SiGe quantum dots. Biaxial strain due to lattice mismatch between Si and SiGe leads to a separation of the two valleys oriented along $[001]$ and $[00\overline{1}]$ from the other four conduction band ground state valleys. The heterostructure potential and alloy disorder finally lift the remaining degeneracies. (c) Cubic unit cell of relaxed bulk Si (diamond crystal) composed of two interpenetrating fcc sub-lattices separated by a non-primitive translation along a quarter of the face diagonal. Atoms of the first and second fcc sub-lattice are shown in red and blue, respectively. (d) Biaxial strain along $[100]$ and $[010]$ yields a tetragonal crystal with reduced symmetry. Like in the relaxed crystal, the two sub-lattices are interchangeable by a nonsymmorphic screw symmetry [30, 32]. (e) Additional shear strain along $[110]$ further reduces the symmetry to an orthorhombic system. The displacement between the two interpenetrating sub-lattices is controlled by Kleinman’s internal ionic displacement parameter $\zeta$ (blue arrows). The nonsymmorphic screw symmetry, which maps the two sub-lattices onto each other, is broken by the shear strain.
@@
Recently, several heuristic strategies have been proposed to enhance the valley splitting by engineering the Si/SiGe heterostructure [33, 30, 18, 32]. These include sharp interfaces, narrow QWs and QWs with uniform low Ge-concentration [17] and more elaborate epitaxial profiles in the QW such as a Ge-spike [34] or an oscillating Ge concentration, known as the wiggle well [33, 30, 32]. In all these concepts, there is a complex interplay of resonances, strain and disorder effects, which can lead to an enhancement of the expected valley splitting. For practical applications, a distinction between deterministic and disorder-induced enhancements is crucial to assess the likelihood for the occurrence of spin-valley hotspots.
@@
In order to improve the understanding of the complex physics determining the valley splitting in Si/SiGe qubits, it is pertinent to develop a comprehensive theoretical model that combines all of the aforementioned effects – namely strain, random alloy disorder and non-trivial resonances – into a unified framework. This objective has been accomplished in the present paper by means of an effective mass model, which merges several existing concepts consistently in a unified framework. We demonstrate that our model faithfully reproduces several known results on valley splitting statistics, interface effects and the shear strain-dependency of the long-period wiggle well, but also extends the state of the art.
The paper is organized as follows: In Sec. II we present the theoretical model for the valley splitting in Si/SiGe quantum dots, which involves perturbation theory for the multi-valley coupled envelope wave function model, a model for the random alloy disorder in the heterostructure and the empirical pseudopotential method to account for the electronic band structure and crystal symmetries. Numerical results are described in Sec. III. Major attention is devoted to strain-induced effects, which extend previous findings on the QW interface-width dependency and wiggle well-type heterostructures. Finally, Sec. IV provides a thorough discussion of the results and comparison with similar models followed by an outlook in Sec. V. Several technical considerations on the derivation of the multi-valley coupled envelope equation model, band structure coefficients and valley splitting statistics can be found in the appendix.
@@ II.1 Coupled Envelope Equations
$H_{0}\left(\mathbf{r}\right)=-\frac{\hbar^{2}}{2m_{t}}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)-\frac{\hbar^{2}}{2m_{l}}\frac{\partial^{2}}{\partial z^{2}}+U\left(\mathbf{r}\right)$ (2)
and
@@
where $m_{l}$ and $m_{t}$ are the effective mass tensor components at the silicon conduction band minimum and $U\left(\mathbf{r}\right)$ is the total confinement potential. The intervalley coupling is described by $V_{c}\left(\mathbf{r}\right)$, which involves the plane wave expansion coefficients $c_{\pm}\left(\mathbf{G}\right)=c_{\pm\mathbf{k}_{0}}\left(\mathbf{G}\right)$ of the lattice-periodic part of the Bloch factors at the two valleys $u_{\mathbf{\pm k}_{0}}\left(\mathbf{r}\right)=\sum_{\mathbf{G}}\mathrm{e}^{i\mathbf{G}\cdot\mathbf{r}}c_{\mathbf{\pm k}_{0}}\left(\mathbf{G}\right)$. The band index is suppressed throughout this paper, as we are solely concerned with the (lowest-energy) conduction band. A detailed derivation of the coupled envelope equation model (1) can be found in Appendix A.
@@
where $\omega_{x}$ and $\omega_{y}$ describe the lateral extension of the QD an thereby the orbital splitting $\Delta E_{\mathrm{orb}}=\min\left(\hbar\omega_{x},\hbar\omega_{y}\right)$. In the limiting case of $\omega_{x}=\omega_{y}$, the QD takes a circular shape. Finally, we assume a constant electric field $F$ along the growth direction, which induces the corresponding potential
@@ II.2 Heterostructure Potential and Alloy Disorder
The simple covariance function reflects the assumption of locally independent distribution of Ge atoms stated in Eq. (8) (i.e., no clustering of Ge atoms). The covariance function (14) determines the statistical properties of the intervalley-coupling parameter and the valley splitting, see Appendix C for details.
@@ II.3 Empirical Pseudopotential Theory and Strain
The empirical pseudopotential method (EPM) provides an accurate description of the electronic band structure with only a few parameters adapted to experimental data. The approach is particularly efficient for silicon, germanium and other semiconductors [46, 47]. A major advantage of the EPM is that it can naturally account for strain effects arising from a displacement of the crystal ions $\mathbf{R}_{i}^{\prime}=\left(I+\varepsilon\right)\mathbf{R}_{i}$, where $\varepsilon$ is the strain tensor. The key steps for the inclusion of strain in EPM calculations are the consideration of strained reciprocal lattice vectors $\mathbf{G}_{i}^{\prime}\approx\left(I-\varepsilon\right)\mathbf{G}_{i}$ (assuming small strain to linear order), the modification of the primitive unit cell volume $\Omega_{p}^{\prime}\approx\left(1+\mathrm{tr}\left(\varepsilon\right)\right)\Omega_{p}$, interpolation of the atomic pseudopotentials at the strained lattice vectors and consideration of internal ionic displacement, describing the relative shift between the two sub-lattices, see Fig. 1 (e). For strained SiGe alloys, numerous empirical pseudopotential models can be found in the literature [48, 49, 50, 51, 52, 42, 53].
In this paper, we use the non-local EPM model described by Fischetti & Laux [49]. From this model, we have obtained the plane wave expansion coefficients of the Bloch factors $c_{\mathbf{k}}\left(\mathbf{G}\right)$, the location of the conduction band minima ±$\pm\mathbf{k}_{0}=\left(0,0,\pm k_{0}\right)$ (i.e., the two valley states) and the corresponding effective mass tensor components $m_{l}$ and $m_{t}$ by numerical diagonalization of the pseudopotential Hamiltonian (taking 113 plane waves into account). We assume biaxial (tensile) strain due to the lattice mismatch between the $\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ substrate and the Si quantum well [54]
@@ II.4 Perturbation Theory
The valley splitting can be approximated by first-order degenerate perturbation theory [14] assuming that both the intervalley-coupling potential $V_{c}\left(\mathbf{r}\right)$ and the alloy-disorder potential $\delta U_{\mathrm{\mathrm{het}}}\left(\mathbf{r}\right)$ are small. In this case, the problem (1) is to zeroth-order approximated by
@@
In first-order perturbation theory, the multi-valley envelope function in Eq. (1) for the degenerate ground state is taken as $\left(\Psi_{+}\left(\mathbf{r}\right),\Psi_{-}\left(\mathbf{r}\right)\right)^{T}\approx\left(\alpha_{+},\alpha_{-}\right)^{T}\Psi_{0}\left(\mathbf{r}\right)$, which after integration with $\Psi_{0}^{*}\left(\mathbf{r}\right)$ yields the reduced problem
$\left(\begin{array}[]{cc}E_{0}+\delta E_{0}&\Delta\\ \Delta^{*}&E_{0}+\delta E_{0}\end{array}\right)\left(\begin{array}[]{c}\alpha_{+}\\ \alpha_{-}\end{array}\right)=E\left(\begin{array}[]{c}\alpha_{+}\\ \alpha_{-}\end{array}\right).$ (19)
@@
is the complex-valued intervalley-coupling strength. We note that Eq. (20) involves all of the aforementioned effects, i.e., strain (via modification of the reciprocal lattice vectors and the Bloch factor expansion coefficients), alloy disorder and non-trivial resonances due to coupling with valley states in neighboring Brillouin zones. Finally, the energy splitting of the two valley states described by the reduced problem (19) is obtained as
@@ II.5.1 Deterministic Component
Deterministic enhancements of the valley splitting can be achieved, when the product of the confinement potential and the longitudinal wave function component resonate with the complex exponential in Eq. (24) [30]. This means, that the Fourier spectrum of
@@
gives a strong contribution via an enhancement of the coupling strength between certain valley states, possibly across different Brillouin zones. Here we used $S\left(z\right)=\sum_{q}\mathrm{e}^{iqz}\tilde{S}\left(q\right)$ and $\tilde{S}\left(-q\right)=\tilde{S}^{*}\left(q\right)$. A similar reasoning holds true for an enhancement of the transverse component (25) with correspondingly adapted $S\left(z\right)$, but for typical parameter values the longitudinal component (24) is dominant. This concept of confinement potential engineering to achieve large Fourier amplitudes $\tilde{S}\left(q\right)$ is explicitly addressed in the wiggle well proposal [33, 30, 32], but is implicitly exploited also in other approaches (e.g., sharp interfaces [23], Ge-spike [34]).
@@
where we have introduced the effective band structure coefficients $C_{n}^{\left(2\right)}$ described in Eq. (58) and the integrals $J_{n}^{\mathrm{det}}$ defined in Eq. (68). The compact form allows for a fast numerical evaluation, where a single summation over a few integers replaces the tedious double summation over the reciprocal lattice vectors in Eq. (24)–(25).
Fig. 2: (a) Band structure coefficients $C_{n}^{\left(2\right)}$ governing the magnitude of the deterministic contribution to the valley splitting as a function of shear strain $\varepsilon_{x,y}$. The coefficients have been computed from Eq. (58) using plane wave expansions of the Bloch factors for strained silicon at the conduction band minimum. Next to shear strain, biaxial tensile strain arising from the $\mathrm{Si}/\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ heterostructure was assumed, see Eq. (15). Coefficients with odd $n$ show a linear dependency on shear strain, while coefficients with even $n$ are practically constant. (b) Same as (a) for the band structure coefficients $C_{n}^{\left(4\right)}$ given by Eq. (67) governing the magnitude of the disorder-induced contribution to the valley splitting.
The above expressions fully account for strain (to linear order) and non-trivial resonance due to coupling with valley states in neighboring Brillouin zones. Hence, the present model goes beyond the commonly employed “$2k_{0}$-theory”, which accounts only for the $n=0$ contribution in Eq. (27). This corresponds to the trivial resonance condition $\mathbf{G}=\mathbf{G}^{\prime}$, cf. Ref. [15], where only the spectral component at
@@
contributes. The plot of the band structure coefficients in Fig. 2 (a) shows, that this approximation is indeed justified in many scenarios, but might fail when the integrands of Eqs. (24)–(25) exhibit strong non-trivial resonances, i.e., with $\mathbf{G}-\mathbf{G}^{\prime}=n\mathbf{G}_{0}$ for $n\neq 0$ and $\mathbf{G}_{0}$ given in Eq. (62). In these cases, contributions proportional to band structure coefficients $C^{\left(2\right)}_{n\neq 0}$ become relevant. A particularly important resonance is the one mediated by $C_{n=-1}^{\left(2\right)}$ at
@@
that is targeted in the long-period wiggle-well [33, 30, 32]. Here we introduced
@@
which is the reciprocal-space distance of the valley states from the Brillouin zone boundary (corresponding to the nearest $X$-point in the absence of strain), see Fig. 1 (a). We note that uniaxial strain $\varepsilon_{z,z}$ (which is typically compressive, see Eq. (15)) leads to a slight modification of this distance. As shown in Fig. 2, the band structure coefficient $C_{n=-1}^{\left(2\right)}$ obtained from the EPM is linearly dependent on the shear strain component as $C_{n=-1}^{\left(2\right)}\approx-0.589\times\varepsilon_{x,y}$. This result is in good agreement with the theory recently presented by Woods et al. [32], which predicts a linear dependency of the long-period wiggle-well on shear strain following a $\mathrm{s}\mathrm{p}^{3}\mathrm{d}^{5}\mathrm{s}^{*}$ tight-binding model. One can clearly see that the resonance is suppressed in the absence of shear strain $\varepsilon_{x,y}=0$, which is explained by a nonsymmorphic screw symmetry of the (relaxed or biaxially strained) silicon crystal structure [30, 32]. In the presence of shear strain, however, this symmetry is broken such that the $n=-1$ resonance yields enhancements also for other strategies much simpler than the long-period wiggle-well, e.g., sharp interfaces or uniform Ge concentrations in the QW. This is discussed in more detail along with numerical simulation results in Sec. III below. Finally, in Sec. III we also observe small contributions of the non-trivial $n=-2$ resonance, in particular at sharp interfaces. This corresponds to
@@
i.e., a coupling between distant valley states separated by an intermediate Brillouin zone. The impact of further non-trivial resonance was found to be negligible.

### v2

@@ Abstract
Electron spin-qubits in silicon-germanium (SiGe) heterostructures are a major candidate for the realization of scalable quantum computers. A critical challenge in strained Si/SiGe quantum wells (QWs) is the existence of two nearly degenerate valley states at the conduction band minimum that can lead to leakage of quantum information. To address this issue, various strategies have been explored to enhance the valley splitting (i.e., the energy gap between the two low-energy conduction band minima), such as sharp interfaces, oscillating germanium concentrations in the QW (known as wiggle well) and shear strain engineering. In this work, we develop a comprehensive envelope-function theory augmented by empirical nonlocal pseudopotential theory to incorporate the effects of alloy disorder, strain, and non-trivial resonances arising from interaction between valley states across different Brillouin zones. We apply our model to analyze common epitaxial profiles studied in the literature with a focus on wiggle well type structures and compare our results with previous work. Our framework provides an efficient tool for quantifying the interplay of these effects on the valley splitting, enabling complex epitaxial profile optimization in future work.
@@ I Introduction
Electron spin-qubits in Si/SiGe quantum dots (QDs) are one of the major candidates for the realization of fault-tolerant universal quantum computers [1, 2, 3]. The material platform has excellent scalability prospects because of the abundance of nuclear spin free isotopes (e.g., 28Si and 76Ge) required for long coherence times and its compatibility with industrial fabrication technology [4, 5, 6]. Experiments have demonstrated high-fidelity state initialization and readout in combination with one and two-qubit gates exceeding the fault-tolerance threshold [7, 8, 9]. Scalable quantum computing architectures require coherent coupling of distant qubits to overcome crosstalk and qubit wiring limitations [10, 11]. As a major step in this direction, coherent qubit transfer across the chip was recently demonstrated using conveyor-mode electron spin-qubit shuttles [12, 13, 14].
@@
One of the key challenges in the design of reliable Si/SiGe qubits is the enhancement of the energy splitting between the two nearly degenerate valley states at the conduction band minimum of a biaxially strained SiGe/Si/SiGe quantum well (QW), see Fig. 1. The energy splitting between these states, called valley splitting, is caused by the coupling of the two valley states by the heterostructure potential [15, 16, 17]. Interface roughness and random alloy disorder in the SiGe barrier [18, 19, 20] cause the valley splitting to fluctuate across the chip with typical values ranging from several tens to hundreds of µeV [21, 22, 23, 24]. These statistical fluctuations of the valley splitting are notoriously difficult to control and inevitably lead to spatial domains with very low splitting, where it becomes comparable with the Zeeman-splitting [25, 26, 23]. As a result, low valley splittings lead to so-called spin-valley hotspots, which are a potential source for spin-dephasing and leakage of quantum information [27, 28]. While these issues might be secondary for stationary qubits, they are particularly critical in spin-qubit shuttles, where the electron is conveyed over micrometer distances across a disordered landscape [29, 30]. For such applications a reliably large valley splitting above 200 µeV would be desirable to avoid spin-valley hotspots [31, 29, 32].
@@
Fig. 1: (a) First Brillouin zone of the face-centered cubic (fcc) lattice. The degeneracy between the six equivalent conduction band minima in Si near the $X$-points is lifted by strain. (b) Energy diagram for the conduction band ground states in Si/SiGe quantum dots. Biaxial strain due to lattice mismatch between Si and SiGe leads to a separation of the two valleys oriented along $[001]$ and $[00\overline{1}]$ from the other four conduction band ground state valleys. The heterostructure potential and alloy disorder finally lift the remaining degeneracies. (c) Cubic unit cell of relaxed bulk Si (diamond crystal) composed of two interpenetrating fcc sub-lattices separated by a non-primitive translation along a quarter of the face diagonal. Atoms of the first and second fcc sub-lattice are shown in red and blue, respectively. (d) Biaxial strain along $[100]$ and $[010]$ yields a tetragonal crystal with reduced symmetry. Similar to the relaxed crystal, the two sub-lattices are interchangeable by a nonsymmorphic screw symmetry [31, 33]. (e) Additional shear strain along $[110]$ further reduces the symmetry to an orthorhombic system. The displacement between the two sub-lattices is controlled by Kleinman’s internal ionic displacement parameter $\zeta$ (blue arrows). The nonsymmorphic screw symmetry, which maps the two sub-lattices onto each other, is broken by the shear strain.
@@
Recently, several heuristic strategies have been proposed to enhance the valley splitting by engineering the Si/SiGe heterostructure [34, 31, 19, 33]. These include sharp interfaces, narrow QWs and QWs with uniform low Ge-concentration [18] and more elaborate epitaxial profiles in the QW such as a Ge-spike [35] or an oscillating Ge concentration, known as the wiggle well [34, 31, 33]. In all these concepts, there is a complex interplay of resonances, strain and disorder effects, which can lead to an enhancement of the mean valley splitting. For practical applications, however, it is important to assess the magnitude of the deterministic enhancement along with the strength of the disorder-induced fluctuations.
@@
In order to improve the understanding of the complex physics determining the valley splitting in Si/SiGe qubits, it is pertinent to develop a comprehensive theoretical model that combines all of the aforementioned effects – namely strain, random alloy disorder and non-trivial resonances – into a unified framework. This objective has been accomplished in the present paper by means of an effective mass model, which merges several existing concepts consistently in a common framework. We demonstrate that our model faithfully reproduces several known results on valley splitting statistics, interface effects and the shear strain-dependency of the long-period wiggle well, but also extends the state of the art.
The paper is organized as follows: In Sec. II , we provide the theoretical model for the valley-splitting in Si/SiGe QDs which involves a multi-valley coupled envelope equation, a statistical model of the random alloy disorder and the emperical pseudopotential method to account for the electronic band structure, strain and crystal symmetries. We provide expressions for the statistical properties of the valley splitting. Numerical results are described in Sec. III. Major attention is devoted to strain-induced effects, which extend previous findings on the QW interface-width dependency and wiggle well-type heterostructures. Finally, Sec. IV provides a thorough discussion of the results and a comparison with similar models followed by an outlook in Sec. V. Several technical considerations on the derivation of the multi-valley coupled envelope equation model, band structure coefficients and valley splitting statistics can be found in the appendix.
@@ II.1 Coupled Envelope Equations
$H_{0}\left(\mathbf{r}\right)=-\frac{\hbar^{2}}{2m_{t}}\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}\right)-\frac{\hbar^{2}}{2m_{l}}\frac{\partial^{2}}{\partial z^{2}}+U\left(\mathbf{r}\right),$ (2)
where $m_{l}$ and $m_{t}$ are the effective mass tensor components at the silicon conduction band minimum and $U\left(\mathbf{r}\right)$ is the total confinement potential. The intervalley coupling is described by
@@
which involves the plane wave expansion coefficients $c_{\pm}\left(\mathbf{G}\right)=c_{\pm\mathbf{k}_{0}}\left(\mathbf{G}\right)$ of the lattice-periodic part of the Bloch factors at the two valleys $u_{\mathbf{\pm k}_{0}}\left(\mathbf{r}\right)=\sum_{\mathbf{G}}\mathrm{e}^{i\mathbf{G}\cdot\mathbf{r}}c_{\mathbf{\pm k}_{0}}\left(\mathbf{G}\right)$. The band index is suppressed throughout this paper, as we are solely concerned with the (lowest-energy) conduction band. A detailed derivation of the coupled envelope equation model (1) can be found in Appendix A.
@@
where $\omega_{x}$ and $\omega_{y}$ describe the lateral extension of the QD an thus the orbital splitting $\Delta E_{\mathrm{orb}}=\min\left(\hbar\omega_{x},\hbar\omega_{y}\right)$. In the limiting case of $\omega_{x}=\omega_{y}$, the QD takes a circular shape. Finally, we assume a constant electric field $F$ along the growth direction, which induces the potential
@@ II.2 Heterostructure Potential and Alloy Disorder
The simple covariance function reflects the assumption of locally independent distribution of Ge atoms stated in Eq. (8) (i.e., no clustering of Ge atoms) and determines the statistical properties of the intervalley-coupling parameter, see Appendix C for details.
@@ II.3 Empirical Pseudopotential Theory and Strain
The empirical pseudopotential method (EPM) provides an accurate description of the electronic band structure with only a few parameters fitted to experimental data. The approach is particularly efficient for silicon, germanium and other semiconductors [47, 48, 49]. A major advantage of the EPM is that it can naturally account for strain effects arising from a displacement of the crystal ions $\mathbf{R}_{i}^{\prime}=\left(I+\varepsilon\right)\mathbf{R}_{i}$, where $\varepsilon$ is the strain tensor. The key steps for the inclusion of strain in EPM calculations are the consideration of strained reciprocal lattice vectors $\mathbf{G}_{i}^{\prime}\approx\left(I-\varepsilon\right)\mathbf{G}_{i}$ (assuming small strain to linear order), the modification of the primitive unit cell volume $\Omega_{p}^{\prime}\approx\left(1+\mathrm{tr}\left(\varepsilon\right)\right)\Omega_{p}$, interpolation of the atomic pseudopotentials at the strained lattice vectors and consideration of internal ionic displacement, describing the relative shift between the two sub-lattices, see Fig. 1 (e). For strained SiGe alloys, numerous empirical pseudopotential models can be found in the literature [50, 51, 52, 53, 54, 43, 55].
In this paper, we use the nonlocal EPM model described by Fischetti & Laux [51]. From this model, we have obtained the plane wave expansion coefficients of the Bloch factors $c_{\mathbf{k}}\left(\mathbf{G}\right)$, the location of the conduction band minima ±$\pm\mathbf{k}_{0}=\left(0,0,\pm k_{0}\right)$ and the corresponding effective mass tensor components $m_{l}$ and $m_{t}$ by numerical diagonalization of the pseudopotential Hamiltonian (taking 113 plane waves into account). We assume biaxial (tensile) strain due to the lattice mismatch between the $\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ substrate and the Si quantum well [56]
@@ II.4 Perturbation Theory
The valley splitting can be approximated by first-order degenerate perturbation theory [15] assuming that both the intervalley-coupling potential $V_{c}\left(\mathbf{r}\right)$ and the alloy-disorder potential $\delta U_{\mathrm{\mathrm{het}}}\left(\mathbf{r}\right)$ are small. In this case, the problem (1) is to zeroth order approximated by
@@
In first-order perturbation theory, the multi-valley envelope function in Eq. (1) for the degenerate ground state subspace is taken as $\left(\Psi_{+}\left(\mathbf{r}\right),\Psi_{-}\left(\mathbf{r}\right)\right)^{T}\approx\left(\eta_{+},\eta_{-}\right)^{T}\Psi_{0}\left(\mathbf{r}\right)$, which after integration with $\Psi_{0}^{*}\left(\mathbf{r}\right)$ yields the reduced problem
$\left(\begin{array}[]{cc}E_{0}+\delta E_{0}&\Delta\\ \Delta^{*}&E_{0}+\delta E_{0}\end{array}\right)\left(\begin{array}[]{c}\eta_{+}\\ \eta_{-}\end{array}\right)=E\left(\begin{array}[]{c}\eta_{+}\\ \eta_{-}\end{array}\right).$ (19)
@@
is the complex-valued intervalley-coupling strength. We note that Eq. (20) involves all of the aforementioned effects, i.e., strain (via modification of the reciprocal lattice vectors and the Bloch factor expansion coefficients), alloy disorder and non-trivial resonances due to coupling of valley states within different Brillouin zones. Finally, the energy splitting of the two valley states described by the reduced problem (19) is obtained as
@@ II.5.1 Deterministic Component
Deterministic enhancements of the valley splitting can be achieved, when the product of the confinement potential and the longitudinal wave function component resonate with the complex exponential in Eq. (24), cf. Ref. [31]. This means, that the Fourier spectrum of
@@
gives a strong contribution via an enhancement of the coupling strength between certain valley states, possibly across different Brillouin zones. A similar reasoning holds true for an enhancement of the transverse component (25) with correspondingly adapted $S\left(z\right)$, but for typical parameter values the longitudinal component (24) is dominant. This concept of confinement potential engineering to achieve large Fourier amplitudes $\tilde{S}\left(q\right)$ is explicitly addressed in the wiggle well proposal [34, 31, 33], but is implicitly exploited also in other approaches (e.g., sharp interfaces [24], Ge-spike [35]).
@@
where we have introduced the effective band structure coefficients $C_{n}^{\left(2\right)}$ described in Eq. (58) and the integrals $J_{n}^{\mathrm{det}}$ defined in Eq. (68). The compact form allows for a fast numerical evaluation, where a single summation over a few integers replaces the tedious double summation over the reciprocal lattice vectors in Eqs. (24)–(25).
Fig. 2: (a) Band structure coefficients $C_{n}^{\left(2\right)}$ governing the magnitude of the deterministic contribution to the valley splitting as a function of shear strain $\varepsilon_{x,y}$. The coefficients have been computed from Eq. (58) using plane wave expansions of the Bloch factors for strained silicon at the conduction band minimum. In addition to shear strain, biaxial tensile strain arising from the $\mathrm{Si}/\mathrm{Si}_{0.7}\mathrm{Ge}_{0.3}$ heterostructure was assumed, see Eq. (15). Coefficients with odd $n$ show a linear dependence on shear strain, while coefficients with even $n$ are practically constant. (b) Same as (a) for the band structure coefficients $C_{n}^{\left(4\right)}$ given by Eq. (67) governing the magnitude of the disorder-induced contribution to the valley splitting.
The above expressions fully account for strain (to linear order) and non-trivial resonances due to coupling with valley states in neighboring Brillouin zones. Hence, the present model goes beyond the commonly employed “$2k_{0}$-theory”, which accounts only for the $n=0$ contribution in Eq. (27). This corresponds to the trivial resonance condition $\mathbf{G}=\mathbf{G}^{\prime}$, cf. Ref. [16], where only the spectral component at
@@
is taken into account. The plot of the band structure coefficients in Fig. 2 (a) shows, that this approximation is indeed justified in many scenarios, but might fail when the integrands of Eqs. (24)–(25) exhibit strong non-trivial resonances, i.e., with $\mathbf{G}-\mathbf{G}^{\prime}=n\mathbf{G}_{0}$ for $n\neq 0$ and $\mathbf{G}_{0}$ given in Eq. (62). In these cases, contributions proportional to band structure coefficients $C^{\left(2\right)}_{n\neq 0}$ become relevant. A particularly important resonance is the one mediated by $C_{n=-1}^{\left(2\right)}$ at
@@
that is associated with the long-period wiggle-well [34, 31, 33]. Here we introduced
@@
which is the reciprocal-space distance of the valley states from the Brillouin zone boundary, see Fig. 1 (a). We note that uniaxial strain $\varepsilon_{z,z}$ (which is typically compressive) leads to a slight modification of this distance. As shown in Fig. 2, the band structure coefficient $C_{n=-1}^{\left(2\right)}$ obtained from the EPM is linearly dependent on the shear strain component as $C_{n=-1}^{\left(2\right)}\approx-0.589\times\varepsilon_{x,y}$. This result is in good agreement with the theory recently presented by Woods et al. [33], which predicts a linear dependency of the long-period wiggle-well on shear strain following a $\mathrm{s}\mathrm{p}^{3}\mathrm{d}^{5}\mathrm{s}^{*}$ tight-binding model. One can clearly see that the resonance is suppressed in the absence of shear strain $\varepsilon_{x,y}=0$, which is explained by a nonsymmorphic screw symmetry of the (either relaxed or biaxially strained) silicon crystal structure [31, 33]. In the presence of shear strain, however, this symmetry is broken such that the resonance at $n=-1$ yields enhancements also for other strategies much simpler than the long-period wiggle-well, e.g., sharp interfaces or uniform Ge concentrations in the QW. This is discussed in more detail along with numerical results in Sec. III below. Finally, in Sec. III we also observe small contributions from the non-trivial $n=-2$ resonance, in particular at sharp interfaces. This corresponds to
@@
i.e., a coupling between distant valley states separated by an intermediate Brillouin zone. The impact of further non-trivial resonances was found to be negligible.